OpenStudy (aye_my_peeps):

dan had $160 in his account. After depositing his paycheck,he had $224 in his bank account. what was the percent of increase in dan bank account after the deposit?

563blackghost (563blackghost):

\(\LARGE{Percent~Increase=\frac{ New~Amount-Old~Amount}{ Old~Amount } \times 100}\)

OpenStudy (aye_my_peeps):

@xXAnnTDMXx

OpenStudy (seratul):

Hey @AYE_MY_PEEPS Can you tell me how much more money Dan has now from before?

563blackghost (563blackghost):

The new amount would be after the deposit and the old amount would be before the deposit...

563blackghost (563blackghost):

\(\LARGE\color{red}{Percent~Increase=\frac{ 224-160}{ 160} \times 100} \)

OpenStudy (aye_my_peeps):

0.4?

563blackghost (563blackghost):

That would be correct but we need to multiply that by 100 to get the percent... \(\Huge{.4 \times 100 = ? }\)

OpenStudy (aye_my_peeps):

40?

OpenStudy (misssmartiez):

Black ghost, is the formula for percent of decrease: Percent decrease = (Older – Newer) ÷ Older? or increase.

OpenStudy (aye_my_peeps):

is it 40%

OpenStudy (aye_my_peeps):

thnx bye

OpenStudy (misssmartiez):

Yes.

563blackghost (563blackghost):

Thats correct ^^

OpenStudy (misssmartiez):

Hey, is it?

563blackghost (563blackghost):

Percent Increase and Percent Decrease equations....

OpenStudy (misssmartiez):

Ah! So that is what I was doing wrong. I was doing determine the missing percent. Not increase and decrease.

OpenStudy (misssmartiez):

Thanks c:

563blackghost (563blackghost):

Np ^^